OpenStudy (anonymous):

I cannot figure out the first step I take in this equation 5t+2(t-1)/7=4 Help please

OpenStudy (anonymous):

Expand 2(t-1) and then multiply everything by seven to get rid of the fraction

OpenStudy (anonymous):

35t+2t-2=28 I feel like thats not right?

OpenStudy (anonymous):

5t+(2t-2)/7=4 35t+2t-2=28 37t=30 t=30/37

OpenStudy (anonymous):

\[2d-1/2(2d-4)=-5/4(d+4)\] Ok on this one would I distribute the fractions or multiply by the fractions inverse tof them?

OpenStudy (anonymous):

*to get rid of them?